Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March):

Molding Fabrication Total
Estimated total machine-hours used 2,500 1,500 4,000
Estimated total fixed manufacturing overhead $ 14,250 $ 17,550 $ 31,800
Estimated variable manufacturing overhead per machine-hour $ 3.10 $ 3.90
Job P Job Q
Direct materials $ 30,000 $ 16,500
Direct labor cost $ 34,600 $ 14,300
Actual machine-hours used:
Molding 3,400 2,500
Fabrication 2,300 2,600
Total 5,700 5,100

Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month.

Required:

For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments.

10. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do not round intermediate calculations.)
Answer #1

Calculate overhead rate :

Molding department = (14250/2500)+3.10 = $8.80 per machine hour

Fabrication department = (17550/1500)+3.90 = $15.60 per machine hour

10) Overhead applied

Job P = (3400*8.80+2300*15.60) = $65800

Job Q = (2500*8.80+2600*15.60) = $62560